OpenStudy (shaylasha):

Simplify (4xy)(2x2y)(3xy)3.

OpenStudy (harman.singh):

4xy * 2x * 2y * 3x y *3 Group the coefficients and the variables together = 4*2*2*3*3*x*x*x*y*y*y Do you know where to go from here?

OpenStudy (harman.singh):

Rule for multiplying powers \[x^a * x^b = x ^{\left( a+b \right)}\]

OpenStudy (triciaal):

just checking, are the parenthesis in the correct places? @harman.singh please check your expression

OpenStudy (harman.singh):

Hey @triciaal , could you please explain what exactly you are referring to? Thanks

OpenStudy (triciaal):